### v1

# LATTE: Learning Aligned Transactions and Textual Embeddings for Bank Clients

Abstract
Learning clients embeddings from sequences of their historic communications is central to financial applications. While large language models (LLMs) offer general world knowledge, their direct use on long event sequences is computationally expensive and impractical in real-world pipelines. In this paper, we propose LATTE, a contrastive learning framework that aligns raw event embeddings with semantic embeddings from frozen LLMs. Behavioral features are summarized into short prompts, embedded by the LLM, and used as supervision via contrastive loss. The proposed approach significantly reduces inference cost and input size compared to conventional processing of complete sequence by LLM. We experimentally show that our method outperforms state-of-the-art techniques for learning event sequence representations on real-world financial datasets while remaining deployable in latency-sensitive environments.
1 Introduction
Learning representations from structured event sequences [26, 27, 28, 29, 30] remains a core challenge in industrial applications of healthcare [36], education [12], e-commerce [10, 37], and, especially, finance [31, 35]. In the banking sector, hundreds of streams of temporally ordered, high-dimensional, and often sparse tabular data [41], such as transaction logs, payment histories, and customer interactions, are generated continuously. These data underpin a broad spectrum of business-critical tasks, including churn prediction, risk assessment, credit scoring, and personalized targeting [13]. Despite abundant customer interaction data, high-quality labels for such downstream tasks remain limited. This shortage of timely supervision hinders the scalability of supervised learning in production settings. It highlights the need for self-supervised approaches to derive robust and semantically rich representations directly from raw behavioral sequences [14]. Prevailing self-supervised approaches for modeling event sequences adopt such objectives as contrastive learning [1], next-event prediction [22], and latent sequence modeling techniques, e.g., Contrastive Predictive Coding (CPC) [23], aiming to capture temporal dependencies and user intent without relying on manual supervision.
Figure 1: Figure of Merit comparing ROC-AUC performance and inference speed (samples/sec/GPU) on the Gender prediction task. Compared methods include LATTE [1], LATTE-S[1], CoLES [1] RTD [24], CPC [23], NSP [25], NPPR [22], DeTPP [33], IFTPP [34], IFTPP-T [34], ESQA [15], and TALLRec [19].
Recent advancements in Large Language Models (LLMs) offer new opportunities to enhance representation learning from event sequences. Trained on diverse and large-scale corpora, LLMs encode elements often implicit or absent in structured event datasets, e.g., rich semantic priors about behavioral patterns, temporal dynamics, and domain knowledge. Leveraging this external knowledge can significantly improve the quality of user representations, particularly in financial applications [43]. Motivated by this potential, recent studies have explored adaptations of LLMs to structured data. For example, TALLRec [19], LLM-TRSR [20], and HKFR [21] transfer rich text understanding abilities of LLMs to recommender systems; TabLLM [18] targets tabular classification tasks; TEST [17] and Time-LLM [16] address time series; while ESQA [15] applies LLMs to event-sequence question answering.
However, a naive direct application of LLMs to serialized sequential tabular data incurs substantial computational overhead due to the large token counts per user. For example, typical banking transaction datasets often contain hundreds of records per user, each serialized into dozens of tokens, exceeding practical context window limits and increasing inference and training times. Figure 1 shows that LLM-based models such as TALLRec [19] and ESQA [15] do not exceed an inference speed of 10 users per second, which severely limits their applicability in banking production environments.
Existing methods to mitigate this issue fall into two main categories. The first reduces context length by summarizing user histories with general-purpose LLMs [21, 20], which risks losing domain-specific information critical for accurate modeling. The second class of methods attempts to bypass context length constraints by encoding sequences in non-textual formats [17, 16]. However, these representations often discard the semantic content present in item descriptions (e.g., transaction categories or merchant details). Moreover, as user histories grow longer, these models either incur increased computational cost or suffer performance degradation due to limited model capacity.
Figure 2: Overview of the proposed LATTE pipeline. (A) Event sequences serve as the input data source. (B) A summary prompt is crafted to query the event sequence. (C) An LLM generator produces a natural language description based on the prompt. (D) The resulting textual description captures salient features of the event sequence. (E) A text embedder converts the description into a vector representation. (F) This description embedding encodes the generated text. (G) In parallel, a sequence encoder embeds the original event sequence. (H) The resulting sequence embedding captures structural and temporal patterns. (I) A contrastive alignment module trains the model to align textual and sequence embeddings in a shared representation space. (J) The aligned embeddings can be used for various downstream tasks such as classification, retrieval, or prediction.
To address these limitations, we propose LATTE  [^1] The source code is available at https://anonymous.4open.science/r/LATTE-3BD2/, a scalable framework for Learning Aligned Transactions and Textual Embeddings. Instead of feeding entire sequences into LLMs, we extract compact client-level statistics and use an instruction-tuned LLM to generate natural language summaries. These summaries serve as weak labels, aligned with pretrained sequence embeddings from a lightweight encoder via contrastive learning. At inference, LATTE supports two modes: a standalone encoder (LATTE-S) that retains LLM-level semantics without added overhead, and a combined encoder (LATTE) that fuses textual and structural representations.
To evaluate trade-offs between performance and efficiency, we introduce a Figure of Merit (FOM) comparing ROC-AUC and inference speed on the standard banking dataset of gender prediction task [3] (Figure 1). We evaluate LATTE in two variants, combining two embedding strategies: structural-only (LATTE-S) and concatenated with textual features (LATTE). In terms of performance, all versions of LATTE outperform the typical baseline financial methods. The structural-only variant (LATTE-S) achieves a ROC-AUC of 0.899 on the Gender task with an inference speed of 162 samples/sec/GPU, surpassing LLM-based models such as TALLRec and ESQA, being over $14\times$ faster. Across all types of textual encoders, the combined variant (LATTE) consistently achieves the highest overall ROC-AUC.
2 Proposed Approach
We aim to improve the quality of representations learned from transactional event sequences by introducing an auxiliary modality derived from textual behavioral descriptions. To this end, we propose a three-stage pipeline, LATTE, illustrated in Figure 2, which maps raw transaction sequences into rich embeddings suitable for downstream tasks.
Let $T_{i}={x_{1},x_{2},\dots,x_{n}}$ denote the transaction sequence for client $i$, where each $x_{j}$ contains timestamped categorical and numerical attributes (e.g., amount, merchant category). As shown in Fig. 2A, we first compute a vector of summary features $s_{i}$ that aggregates behavioral patterns over the sequence: frequency of activity, merchant diversity, transaction types, temporal coverage, and income-expense structure. Categorical statistics are transformed into meaningful textual descriptions rather than raw indices. See Appendix C for the prompt template and a sample generated description. This profile is then serialized into a textual prompt (Fig. 2B), which is passed to a frozen instruction-tuned LLM (Fig. 2C) to generate a natural language description $d_{i}$ of the client’s behavior (Fig. 2D). The prompt includes a system message and a structured representation of $s_{i}$ designed to elicit coherent and interpretable responses.
Simultaneously, the raw sequence $T_{i}$ is processed by a GRU-based sequence encoder (Fig. 2G) trained under a self-supervised CoLES objective [1], where each training example consists of two overlapping subsequences (positives) and contrastive negatives from other clients. This produces the sequence embedding $z_{i}^{\text{seq}}$ optimized to capture client-specific behavioral dynamics. In parallel, the description $d_{i}$ is passed through a frozen multilingual sentence encoder (Fig. 2F), producing the text embedding $z_{i}^{\text{text}}$ (Fig. 2E) via mean pooling.
The embeddings $z_{i}^{\text{seq}}$ and $z_{i}^{\text{text}}$ are then aligned using one of three cross-modal contrastive losses (Fig. 2I), while keeping the text encoder fixed. The updated transaction embeddings $\left(z_{i}^{\text{seq}}\right)^{{}^{\prime}}$ (Fig. 2H), aligned with the textual representations, are evaluated on downstream tasks such as churn prediction (Fig. 2J).
To perform cross-modal alignment, we introduce three contrastive heads: Symmetric Softmax, Pairwise Binary, and Orthogonal Regularized. Each was inspired by prior work on multimodal learning [6, 7, 32]. These heads differ in their alignment geometry and regularization, and we refer to them as LATTE [1], LATTE [2], and LATTE [3], respectively. All three heads use frozen text encoders and update only the transaction encoder. Downstream tasks are evaluated using the resulting $\left(z_{i}^{\text{seq}}\right)^{{}^{\prime}}$, while the textual modality provides semantically grounded alignment by enriching the training signal with contextual knowledge of categorical attributes that the sequence encoder alone cannot interpret.
LATTE [1]: Symmetric Softmax Contrastive Head promotes bidirectional alignment between modalities using a symmetric InfoNCE-style loss:
$\mathcal{L}_{\text{softmax}}=\frac{1}{2}(\mathcal{L}_{\text{seq}\rightarrow\text{text}}+\mathcal{L}_{\text{text}\rightarrow\text{seq}}),$ (1)
where
$\mathcal{L}_{\text{seq}\rightarrow\text{text}}$ $=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp(\langle z_{i}^{\text{seq}},z_{i}^{\text{text}}\rangle/\tau)}{\sum_{j=1}^{N}\exp(\langle z_{i}^{\text{seq}},z_{j}^{\text{text}}\rangle/\tau)}.\$
The second term, $\mathcal{L}_{\text{text}\rightarrow\text{seq}}$, is defined similarly with the roles of sequence and text reversed. Here, $\langle\cdot,\cdot\rangle$ denotes the cosine similarity between the L2-normalized embeddings of the sequence ($z_{i}^{\text{seq}}$) and the corresponding text ($z_{i}^{\text{text}}$), and $\tau$ is a temperature hyperparameter controlling the sharpness of the similarity distribution.
LATTE [2]: Pairwise Binary Contrastive Head applies sigmoid-based supervision across all sequence-description pairs in the batch:
$\mathcal{L}_{\text{sigm}}=-\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{N}\log\left(\frac{1}{1+e^{y_{ij}(-\tau z_{i}^{\text{seq}}\cdot z_{j}^{\text{text}}+b)}},\right)$ (2)
where $y_{ij}=1$ for $i=j$ (positive pairs), $y_{ij}=-1$ otherwise (negatives), and $b$ is a trainable bias.
LATTE [3]: Orthogonal Regularized Contrastive Head augments the softmax-based loss with a geometric regularization term that disentangles modality-specific and shared information. To achieve this, we introduce an auxiliary projection head that maps each transaction embedding $z_{i}^{\text{seq}}$ to a representation composed of two parts: $Z^{\text{shared}}$, which captures components aligned with textual information, and $Z^{\text{spec}}$, which preserves information specific to the transaction modality:
$\mathcal{L}_{\text{reg}}=\mathcal{L}_{\text{softmax}}+\lambda_{\text{ortho}}\cdot\mathcal{L}_{\text{ortho}},$ (3)
where $\mathcal{L}_{\text{ortho}}=\left\lVert(Z^{\text{shared}})^{\top}Z^{\text{spec}}\right\rVert_{F}^{2}$.
This separation promotes disentangled features by penalizing correlation between shared and specific subspaces, where $\lambda_{\text{ortho}}$ controls the strength of this regularization and thus the emphasis on preserving modality-specific information.
3 Experimental Setup
This section provides the core details of our experimental setup, including validation strategy, datasets, and baseline methods. Additional implementation details are provided in Appendix A.
3.1 Data
We evaluate our method on three real-world datasets containing anonymized credit card transaction sequences from major financial institutions. Each dataset comprises client-level sequences with numerical and categorical attributes (e.g., amount, merchant category, transaction type), and includes an unlabeled subset used exclusively for representation learning. Churn [2] includes approximately 10K Rosbank clients labeled by future inactivity. Gender [3] and Age Group [4], provided by Sberbank, contain 15K and 50K clients respectively, annotated with demographic labels.
3.2 Validation Strategy
Each dataset is split into disjoint training and test partitions by reserving 10% of the labeled clients for evaluation. The remaining 90% of labeled users, together with all available unlabeled users, are used for training the embedding models. To assess the quality of learned representations, we adopt a 5-fold cross-validation procedure. Specifically, the labeled portion of the training data is divided into five equal-sized folds. For each fold $v$, we: (1) train a LightGBM [38] classifier on embeddings from the remaining four folds, and (2) evaluate it on the held-out test fold, computing a downstream performance metric $M_{v}$. For binary classification tasks (churn, gender), we report ROC-AUC; for multiclass age prediction, we report classification accuracy. The final performance is summarized as $\mu_{M}\pm\sigma_{M}$, where $\mu_{M}$ is the mean and $\sigma_{M}$ is the standard deviation over all five folds.
3.3 Baselines
We compare LATTE against a diverse set of baselines spanning five methodological families:
Event sequence models. CPC [23] learns to predict future representations from past context via a contrastive loss. CoLES [1] improves temporal consistency by aligning overlapping subsequences using InfoNCE. NPPR [22] employs autoregressive training with dual objectives: predicting the next and reconstructing the previous event from masked sequences. All models operate on raw event sequences without external modalities.
Temporal point process models. DeTPP [33] models event timing and types using parametric point processes. IFTPP and IFTPP-T [34] use Transformer and GRU backbones with combined MAE and classification losses.
Natural language processing objectives. RTD [24] randomly replaces 15% of event tokens and predicts the original token; NSP [25] extends BERT’s next sentence prediction to sequences of events.
LLM-based approaches. We adapt TALLRec [19] and HKFR [21] from recommender systems for user embedding extraction by fine-tuning LLMs on serialized user-item histories using next-token prediction. Embeddings are derived via mean pooling over the final layer, following recent best practices [46, 42]. Additionally, we replace older backbones (e.g., LLaMA 7B [49], ChatGLM-6B [48]) with LLaMA 3.2 3B [47] to leverage architectural improvements and efficiency.
Tabular feature aggregation. As a non-sequential baseline, agg aggregates transaction features using summary statistics such as mean, standard deviation, min-max and grouped frequency statistics (for categorical features).
4 Experimental Results
4.1 Main Results
Model Churn (AUC) Age Group (Acc) Gender (AUC)
agg 0.827 ± 0.010 0.629 ± 0.002 0.877 ± 0.004
CPC 0.792 ± 0.015 0.602 ± 0.004 0.851 ± 0.006
RTD 0.771 ± 0.016 0.631 ± 0.006 0.855 ± 0.008
CoLES 0.841 ± 0.005 0.644 ± 0.005 0.882 ± 0.004
NSP 0.828 ± 0.012 0.621 ± 0.005 0.852 ± 0.011
NPPR 0.845 ± 0.003 0.642 ± 0.001 -
DeTPP 0.823 ± 0.002 0.632 ± 0.004 -
IFTPP 0.828 ± 0.004 0.632 ± 0.003 0.863 ± 0.003
IFTPP-T 0.814 ± 0.004 0.620 ± 0.002 0.852 ± 0.005
TALLRec 0.839 $\pm$ 0.003 0.659 ± 0.004 0.875 $\pm$ 0.004
HKFR 0.823 $\pm$ 0.006 - -
LATTE [1] 0.867 ± 0.004 0.666 ± 0.005 0.900 ± 0.005
LATTE [2] 0.868 ± 0.006 0.671 ± 0.005 0.901 ± 0.005
LATTE [3] 0.857 ± 0.003 0.678 ± 0.003 0.902 ± 0.006
Table 1: Performance of client embeddings on downstream tasks. Bold indicates best result; underline indicates second-best; double underline indicates third-best
Table 1 presents the performance of client embeddings on three downstream tasks: churn prediction (AUC), age group classification (accuracy), and gender prediction (AUC). While traditional self-supervised objectives such as CPC, RTD, and NSP lag behind stronger baselines like CoLES and TALLRec, several variants of our proposed LATTE framework demonstrate clear gains across tasks. LATTE [1] achieves the best result on churn prediction (0.868 AUC), while LATTE [3] leads on both age group classification (0.678 accuracy) and gender prediction (0.902 AUC). These improvements indicate that incorporating synthetic textual supervision leads to consistently stronger representations of transactional behavior.
4.2 Runtime Analysis
Figure 3: Figure of Merit comparing model performance in ROC-AUC and inference speed in samples per second per GPU on Churn dataset.
In banking applications, inference speed is critical due to the need to process millions of multi-source user records in a hardware environment with limited access to GPUs. In this section, we investigate an important trade-off between inference resource utilization (compute time, memory utilization) and the performance of the proposed methods compared to baseline architectures.
Figure 3 showcases the performance on the Churn dataset. Generative models (LATTE [1, 2, 3], TALLRec) achieve the highest ROC-AUC scores, exceeding 0.868, but at the cost of extremely low inference speeds—processing only a few samples per second—and high memory consumption, with over 3 billion parameters. In contrast, lightweight LATTE-S variants attain slightly lower ROC-AUC values while operating over 160 samples per second and compact size of just 4.4 million parameters. These models match the efficiency of contrastive baselines, offering a favorable trade-off between accuracy, speed, and memory footprint. As shown in Appendix B (Figure 4), a similar trend occurs on the Age prediction task.
4.3 Ablation Study
Contrastive Fine-tuning with Textual Embedding Alignment.
Method Churn (AUC) Age Group (Acc) Gender (AUC)
Descriptions 0.718 ± 0.004 0.552 ± 0.002 0.692 ± 0.006
CoLES + descriptions 0.853 ± 0.009 0.662 ± 0.005 0.895 ± 0.008
LATTE-S[1] 0.852 ± 0.005 0.657 ± 0.007 0.890 ± 0.008
LATTE [1] 0.867 ± 0.004 0.666 ± 0.005 0.900 ± 0.005
LATTE-S[2] 0.858 ± 0.007 0.662 ± 0.006 0.894 ± 0.009
LATTE [2] 0.868 ± 0.006 0.671 ± 0.005 0.901 ± 0.005
LATTE-S[3] 0.844 ± 0.005 0.663 ± 0.004 0.899 ± 0.001
LATTE [3] 0.857 ± 0.003 0.678 ± 0.003 0.902 ± 0.006
Table 2: Ablation: Impact of contrastive fine-tuning and modality concatenation on downstream task.
Table 2 summarizes an ablation study on contrastive fine-tuning and modality concatenation. Adding LLM-generated descriptions consistently boosts performance across all heads—for instance, LATTE [2] improves from 0.858 to 0.868 AUC on churn and from 0.894 to 0.901 on gender. Contrastive alignment proves essential, as all LATTE variants outperform CoLES + descriptions, which rely only on feature concatenation. LATTE [3] achieves the best results on age (0.678) and gender (0.902), highlighting the impact of regularization. These results show that language grounding via contrastive supervision outperforms simple fusion.
Language Model Choice for Description Generation.
Contrastive Head LLM Generator Churn (AUC) Age Group (Acc) Gender (AUC)
LATTE [1] Gemma 3 4B 0.849 ± 0.005 0.645 ± 0.006 0.888 ± 0.005
LATTE [1] Gemma 3 27B 0.867 ± 0.004 0.666 ± 0.005 0.900 ± 0.005
LATTE [1] Qwen 3 32B 0.859 ± 0.004 0.658 ± 0.005 0.896 ± 0.003
LATTE [2] Gemma 3 4B 0.851 ± 0.005 0.652 ± 0.006 0.889 ± 0.002
LATTE [2] Gemma 3 27B 0.868 ± 0.006 0.671 ± 0.005 0.901 ± 0.005
LATTE [2] Qwen 3 32B 0.862 ± 0.001 0.667 ± 0.004 0.896 ± 0.000
LATTE [3] Gemma 3 4B 0.849 ± 0.005 0.665 ± 0.006 0.890 ± 0.005
LATTE [3] Gemma 3 27B 0.857 ± 0.003 0.678 ± 0.003 0.902 ± 0.006
LATTE [3] Qwen 3 32B 0.862 ± 0.004 0.675 ± 0.002 0.895 ± 0.003
Table 3: Ablation: Effect of language model choice for generating behavioral descriptions.
Table 3 evaluates how the choice of LLM for behavioral description generation affects downstream task performance. Across all contrastive heads, larger instruction-tuned models consistently yield stronger results. In particular, Qwen 3 32B outperforms smaller alternatives on churn prediction across all heads, achieving the best score of 0.862 AUC under LATTE [1]. Meanwhile, Gemma 3 27B provides competitive performance on age and gender classification, producing the top scores under both LATTE [2] (0.671 / 0.901) and LATTE [3] (0.678 / 0.902). In contrast, the smallest model, Gemma 3 4B, lags across all settings. These results emphasize that the generative capacity of the LLM is crucial for high-quality representation learning.
Text Embedding Model Selection.
Contrastive Head Text Encoder Churn (AUC) Age Group (Acc) Gender (AUC)
LATTE [1] mE5-large-instruct 0.860 ± 0.005 0.658 ± 0.008 0.895 ± 0.003
LATTE [1] Qwen3-Embedding-0.6B 0.862 ± 0.004 0.660 ± 0.007 0.897 ± 0.003
LATTE [1] Qwen3-Embedding-8B 0.867 ± 0.004 0.666 ± 0.005 0.900 ± 0.005
LATTE [2] mE5-large-instruct 0.860 ± 0.007 0.664 ± 0.006 0.894 ± 0.004
LATTE [2] Qwen3-Embedding-0.6B 0.863 ± 0.005 0.667 ± 0.005 0.896 ± 0.003
LATTE [2] Qwen3-Embedding-8B 0.868 ± 0.006 0.671 ± 0.005 0.901 ± 0.005
LATTE [3] mE5-large-instruct 0.850 ± 0.004 0.671 ± 0.003 0.897 ± 0.002
LATTE [3] Qwen3-Embedding-0.6B 0.852 ± 0.006 0.673 ± 0.001 0.899 ± 0.002
LATTE [3] Qwen3-Embedding-8B 0.857 ± 0.003 0.678 ± 0.003 0.902 ± 0.006
Table 4: Ablation: Impact of the text embedding model on downstream task performance.
Table 4 investigates the impact of the frozen text encoder used to obtain $z_{i}^{\text{text}}$ on downstream task performance, where the resulting embeddings serve both for alignment and as a standalone source of world knowledge. We compare mE5-large-instruct [44] with two variants of Qwen3 [45]: a lightweight embedding model (0.6B) and a large one (8B). Across all contrastive heads, Qwen3-Embedding-8B consistently delivers the strongest results, outperforming mE5 by up to 1.0–1.3% on age and gender tasks and up to 0.7% on churn. Notably, under LATTE [3], it pushes age classification to 0.678 and gender AUC to 0.902, both state-of-the-art. Appendix B provides additional ablation comparing different description embedding strategies.
Input Format for Prompting.
Contrastive Head Prompt Format Churn (AUC) Age Group (Acc) Gender (AUC)
LATTE [1] Raw serialization 0.854 ± 0.003 0.653 ± 0.007 0.887 ± 0.004
LATTE [1] Stats 0.867 ± 0.004 0.666 ± 0.005 0.900 ± 0.005
LATTE [2] Raw serialization 0.855 ± 0.005 0.656 ± 0.007 0.886 ± 0.005
LATTE [2] Stats 0.868 ± 0.006 0.671 ± 0.005 0.901 ± 0.005
LATTE [3] Raw serialization 0.842 ± 0.007 0.662 ± 0.005 0.889 ± 0.003
LATTE [3] Stats 0.857 ± 0.003 0.678 ± 0.003 0.902 ± 0.006
Table 5: Ablation: Input format for prompting the LLM raw transaction serialization vs. summary statistics.
Table 5 compares two formats for prompting the LLM: raw serialized transaction sequences and structured behavioral summaries. Across all contrastive heads, using statistical summaries consistently outperforms raw serialization. For instance, in the LATTE [3] setup, statistical prompts improve churn ROC-AUC by 1.5% and gender AUC by 1.3%. Appendix B provides additional ablation of the textual description effect compared to raw statistical features.
5 Conclusion
We presented a novel method for contrastive representation learning from event sequences that leverages synthetically generated textual descriptions as a complementary modality. By aligning structured transaction data with natural language summaries produced by a frozen instruction-tuned LLM, the proposed approach introduces textual priors into the embedding space without requiring labeled data or LLM fine-tuning.
LATTE achieves state-of-the-art results across three key open-source banking tasks, with relative improvements of 6.1% in gender prediction, 3.0% in age group classification, and 2.0% in churn prediction compared to the baseline. The proposed LATTE-S is resource-efficient (4.4M parameters, up to 200 samples/sec speed), which is essential for industrial applications where behavioral logs are abundant but supervision is limited.
A particularly promising future direction is to explore richer forms of text-to-sequence alignment, where natural language summaries are more tightly coupled with the underlying event dynamics. This approach could yield embeddings that are not only effective for predictive tasks but also inherently interpretable, as they remain anchored in human-readable descriptions of user behavior.
Limitations
While our approach demonstrates strong empirical performance, it relies on the assumption that LLM-generated textual descriptions accurately reflect behavioral nuances present in structured sequences. Thus, the prompt design and generalization capacity of the frozen LLM limit the quality of supervision. Additionally, the text encoder is not updated during training, which may limit alignment fidelity in domains with distributional shifts. Finally, although our method does not require labels or fine-tuning, it introduces moderate training overhead compared to lightweight contrastive learning methods due to reliance on LLM-based generation at scale.
Appendix A Experimental Details
For all experiments, natural language descriptions were generated using the instruction-tuned Gemma-3-27B-Instruct [39] model. To encode the resulting behavioral descriptions into fixed-length vectors, we employed the Qwen3-Embedding-8B [40] model, a multilingual encoder optimized for semantic retrieval. The transaction encoder was instantiated as a GRU-based model trained under the CoLES self-supervised framework, serving as the base sequence encoder across all contrastive alignment heads. Both training and inference for the full pipeline, including LLM prompting, embedding computation, and contrastive alignment, were performed on a single NVIDIA Tesla A100 80GB GPU.
Appendix B Additional Experiments
Text Embedding vs. Stats Embedding
Contrastive Head Auxiliary Modality Churn (AUC) Age Group (Acc) Gender (AUC)
LATTE [1] Stats (MLP-1L) 0.846 ± 0.006 0.647 ± 0.006 0.886 ± 0.004
LATTE [1] Stats (MLP-2L) 0.847 ± 0.005 0.645 ± 0.007 0.887 ± 0.003
LATTE [1] LLM text 0.867 ± 0.004 0.666 ± 0.005 0.900 ± 0.005
LATTE [2] Stats (MLP-1L) 0.848 ± 0.005 0.651 ± 0.006 0.883 ± 0.005
LATTE [2] Stats (MLP-2L) 0.851 ± 0.003 0.653 ± 0.005 0.884 ± 0.004
LATTE [2] LLM text 0.868 ± 0.006 0.671 ± 0.005 0.901 ± 0.005
LATTE [3] Stats (MLP-1L) 0.842 ± 0.007 0.657 ± 0.004 0.886 ± 0.003
LATTE [3] Stats (MLP-2L) 0.843 ± 0.005 0.656 ± 0.003 0.887 ± 0.003
LATTE [3] LLM text 0.857 ± 0.003 0.678 ± 0.003 0.902 ± 0.006
Table 6: Ablation: Effect of auxiliary modality used for alignment — handcrafted stats vs. LLM-generated text. Bold indicates best; underline second-best per group.
Table 6 compares two auxiliary supervision targets: raw statistical feature vectors (projected through MLPs) versus LLM-generated natural language descriptions. Text-based supervision leads to stronger downstream performance across all tasks for every contrastive head. The advantage is especially pronounced under LATTE [3], where LLM-based supervision improves age classification accuracy to 0.678 and gender AUC to 0.902 (the highest overall). Even the best-performing 2-layer MLP projections underperform by 1–2% in most settings. These results confirm that natural language descriptions offer more informative targets than handcrafted statistics for guiding sequence representation learning.
Impact of Text Embedding Extraction Strategy.
Contrastive Head Method Churn (AUC) Age Group (Acc) Gender (AUC)
LATTE [1] LLM encoder (ours) 0.867 ± 0.004 0.666 ± 0.005 0.900 ± 0.005
LATTE [1] MeanPool (generator) 0.859 ± 0.006 0.659 ± 0.003 0.893 ± 0.001
LATTE [2] LLM encoder (ours) 0.868 ± 0.006 0.671 ± 0.005 0.901 ± 0.005
LATTE [2] MeanPool (generator) 0.862 ± 0.005 0.666 ± 0.002 0.894 ± 0.003
LATTE [3] LLM encoder (ours) 0.857 ± 0.003 0.678 ± 0.003 0.902 ± 0.006
LATTE [3] MeanPool (generator) 0.851 ± 0.007 0.671 ± 0.004 0.894 ± 0.002
Table 7: Ablation: Comparing description embeddings obtained from a dedicated sentence encoder (ours) vs. directly from the generator LLM.
We compare two strategies for obtaining text embeddings from behavioral descriptions. In our default setup, we use a frozen sentence encoder (Qwen3-Embedding-8B) to compute embeddings, separating the generation and encoding processes. As an alternative, we extract the embedding directly from the generator LLM (Gemma-3-27B) by mean-pooling hidden states. Following recent work suggesting that mean pooling over all token embeddings outperforms using the EOS token [46, 42], we average hidden states over the final $k=8$ transformer layers. For instance, employing a dedicated LLM encoder yields up to a 0.8% improvement in both gender and churn AUC over direct pooling from the generator.
Runtime Analysis
Figure 4: Figure of Merit comparison of model performance in ROC-AUC and inference speed in samples per second per GPU on Age dataset.
Figure 4 presents results for the Age prediction task. We observe trends consistent with those in the Churn dataset (Figure 3): generative models attain the highest accuracy (up to 0.678) but exhibit limited inference throughput. In contrast, lightweight LATTE-S variants achieve a favorable trade-off, maintaining competitive performance ( 0.66 ROC-AUC) while delivering significantly higher inference speeds, exceeding 200 samples per second.
Appendix C Text Generation Protocol
We generate natural language descriptions from statistical user summaries by prompting an instruction-tuned LLM. Table 8 illustrates the full prompt used for behavioral description generation, while Table 9 provides an example output produced by the model.
System: You are an expert in financial transaction analysis. Your task is to generate clear, structured, and concise descriptions of user financial behavior based on given transaction data. Use data-driven insights and avoid speculation.
Guidelines: Do not include phrases like “Here’s a financial behavior description for User X” Start directly with behavioral insights Interpret numbers into patterns (e.g., burst spending, routine payments) Highlight spending habits, risk factors, financial consistency Avoid unsupported assumptions; maintain clarity and conciseness
User: Below is a summary of a user’s transaction history:
- Number of transactions: 226
- Active transaction period: 444 days
- Avg transactions per day: 0.51
- Avg transaction interval: 1.97 days
- Top MCCs: Supermarkets, cash withdrawals, telecommunications
- Share of transactions in Top 1 MCC: 0.04
- Avg amount for Top 1 MCC: 2770.47 RUB
- Avg amount for Top 2 MCC: 45969.41 RUB
- Avg amount for Top 3 MCC: 8139.52 RUB
- Share of days with transactions: 0.36
- Total income: 2,706,553.10 RUB
- Total expenses: 3,956,007.97 RUB
- Avg outgoing amount: 17,981.85 RUB
- Avg incoming amount: 451,092.18 RUB
Instructions: Analyze behavioral patterns Identify transaction regularity and category reliance Assess potential risk factors and financial planning traits Write in a structured and engaging way while staying factual
Table 8: Prompt template for behavioral description generation
Output:
This user demonstrates a low-frequency, high-variance spending pattern over a substantial period (444 days). Transactions occur on approximately 36% of days, averaging one every two days, indicating infrequent but potentially substantial financial activity. Spending is heavily weighted towards cash withdrawals from financial institutions (4% of transactions, average 45969.41 RUB), suggesting a preference for cash-based expenses or large withdrawals for specific purposes. Significant spending also occurs at grocery stores/supermarkets (average 8139.52 RUB), representing a consistent, though smaller, portion of overall expenditure. A smaller number of transactions are related to phone services (average 2770.47 RUB). Expenses significantly exceed income (3,956,007.97 RUB vs. 2,706,553.10 RUB), implying reliance on savings, credit, or other income sources not captured in this data. The high average outgoing transaction amount (17,981.85 RUB) coupled with a large average incoming amount (451,092.18 RUB) suggests irregular but substantial income events are occurring, potentially offsetting consistent overspending. The combination of infrequent transactions, high cash withdrawals, and a substantial expense-to-income imbalance points to potentially inconsistent financial management. While not indicative of reckless spending, the reliance on infrequent income and high cash usage warrants further investigation to determine financial stability and potential risk factors. There’s a lack of diversification in spending, with the top three MCCs dominating activity.
Table 9: Example output generated from the LLM